# Asymmetric Errors

### v2

@@ A.8 The Edgeworth expansion
The cumulative equivalent of this $P(x)$, the distribution function, is
@@
where $\Phi^{-1}$ is the inverse of the cumulative Gaussian. This seems to work better than Newton’s method, thanks to the constantly changing gradient. From $z$ one obtains $M=\sigma z+\mu$, and thence ${\sigma^{+}}=\sigma+\mu-M$ and ${\sigma^{-}}=\sigma-\mu+M$.
@@ A.9 The skew normal
For the quantile parameters we need the cumulative distribution, which is not so elegant
@@ A.10 The Johnson system
The combination of Johnson distributions $S_{B}$ (bounded), $S_{U}$ (unbounded), together with their limiting cases (Gaussian and log normal), form a very flexible system of four-parameter densities [23, 24, 25]. It turns out that there is a Johnson distribution for every possible combination of mean, standard deviation, skewness, and kurtosis. It is in fact convenient to think of the set as if it was a single distribution whose density is given by $j(R,\mu,\sigma,s,\kappa)$. (In this parameterization, skewness $s$ and kurtosis $\kappa$ are normalized).
@@
Figure 26: The Johnson distribution: the panel on the right illustrates the Johnson $S_{U}$ density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Equation 48. As expected, this transformation maps $-1$ to $-0.6$, 00 to 00, and $1$ into $1$.
@@ A.11 The log normal
Figure 27: The log normal distribution: the panel on the right illustrates the density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Equation 50.
@@ B.1 The cubic
If ${\sigma^{+}}$ and ${\sigma^{-}}$ differ by less than a factor of 2 this gives curves which behave sensibly in the range $[\hat{a}-{\sigma^{-}},\hat{a}+{\sigma^{+}}]$, but outside that range the cubic term gives an unwanted turning point. (If the asymmetry is large, the curve will have a minimum inside the range in addition to the maximum at $\hat{a}=0$.) Also the curve does not go to $\infty$- ∞ at both large positive and negative $a$. So although the cubic is the obvious extension of the quadratic, it often gives problems.
@@ B.4 The constrained quartic
$(\ln L)^{\prime\prime}(a)=-{1\over 2}(\alpha+\beta a)^{2}\qquad{\rm which\ yields}\qquad\ln L(a)=-{1\over 2}\left({\alpha^{2}a^{2}\over 2}+{\alpha\beta a^{3}\over 3}+{\beta^{2}a^{4}\over 12}\right)$ (56)
@@
$\beta={1\over{\sigma^{+}}{\sigma^{-}}}\sqrt{{12({\sigma^{-}}+{\sigma^{+}})^{2}\pm 24\sqrt{4{\sigma^{+}}{\sigma^{-}}^{3}+4{\sigma^{-}}{\sigma^{+}}^{3}-2{\sigma^{+}}^{4}-2{\sigma^{-}}^{4}}\over 3{\sigma^{-}}^{2}+2{\sigma^{+}}{\sigma^{-}}+3{\sigma^{+}}^{2}}}$ (57)
@@ B.7 The interpolated 7th degree polynomial
$\begin{split}\alpha&=6({\sigma^{-}}-{\sigma^{+}})/\eta\\ \beta&=15({\sigma^{-}}-{\sigma^{+}})^{2}/\eta\\ \gamma&=10({\sigma^{-}}-{\sigma^{+}})({\sigma^{-}}^{2}-4{\sigma^{-}}{\sigma^{+}}+{\sigma^{+}}^{2})/\eta\\ \delta&=-30{\sigma^{-}}{\sigma^{+}}({\sigma^{-}}-{\sigma^{+}})^{2}/\eta\\ \epsilon&=30{\sigma^{-}}^{2}{\sigma^{+}}^{2}({\sigma^{-}}-{\sigma^{+}})/\eta\\ \zeta&=({\sigma^{-}}^{6}+4{\sigma^{-}}^{5}{\sigma^{+}}+6{\sigma^{-}}^{4}{\sigma^{+}}^{2}-6{\sigma^{-}}^{3}{\sigma^{+}}^{3}+6{\sigma^{-}}^{2}{\sigma^{+}}^{4}+4{\sigma^{-}}{\sigma^{+}}^{5}+{\sigma^{+}}^{6})/\eta\end{split}$ (64)
@@
In combination with Equation 55 for $a<-{\sigma^{-}}$ and $a>{\sigma^{+}}$, this model has a very reasonable behavior for large values of $|a|$ but can be somewhat “wavy” on the interval $[-{\sigma^{-}},{\sigma^{+}}]$. It also has a discontinuous third derivative at $a=-{\sigma^{-}}$ and $a={\sigma^{+}}$.
@@ B.9 The double quintic
$\ln L(a)=\begin{cases}P_{2}^{-}(a)&\text{for }a\leq-{\sigma^{-}}\\ P_{5}^{-}(a)&\text{for }-{\sigma^{-}}\leq a\leq 0\\ P_{5}^{+}(a)&\text{for }0\leq a\leq{\sigma^{+}}\\ P_{2}^{+}(a)&\text{for }{\sigma^{+}}\leq a\end{cases}$ (68)
@@
The simple double quintic is obtained by setting $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$. For this choice of $\sigma_{0}$, the requirement that $\ln L(a)$ has a single extremum (the maximum at 0) leads to the condition that ${\sigma^{+}}$ and ${\sigma^{-}}$ must differ by less than a factor of $13.5$.
@@ B.11 The log logistic-beta
Here, the logarithm of the logistic-beta density (with the $x\rightarrow a$ replacement) has been reparameterized in terms of the new parameters $g>0$, $c>0$, and $|A|<1$, and a constant term has been added so that $\ln L(0)=0$. The role of the new parameters can be appreciated as follows:
@@
Equation 72 satisfies the conditions $\ln L(0)=0$ and $\left.\frac{d\ln L(a)}{da}\right|_{a=0}=0$ automatically. It also needs to satisfy the conditions $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})=-\frac{1}{2}$. While with three parameters $A$, $c$, and $g$ the latter conditions can be satisfied in a multitude of ways, we choose the parameterization that minimizes $g$, i.e. $A$ and $c$ are chosen in such a way that $|\ln L({\sigma^{+}})/g|$ is maximized subject to the constraint $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})$. This results in a conservative log-likelihood model with limited slopes and the second derivative exceeding $\frac{1}{\min({\sigma^{-}},{\sigma^{+}})^{2}}$ in magnitude only on a short interval.
@@ B.13 The generalised Poisson
This has to be solved numerically for $\gamma$. The solution lies between 0 and $1/{\sigma^{-}}$, and can be found by repeatedly taking the midpoint.  [^4] Attempts to use more sophisticated algorithms were unsuccessful. ${\cal N}$ is then found from
@@
This does fairly well, though the need for a numerical solution makes it inelegant. If the curve to be fitted has a negative skewness (${\sigma^{-}}>{\sigma^{+}})$ then the sign of $a$ has to be flipped.
@@ B.14 The linear sigma
$\ln L(a)=-{1\over 2}\left({a-\hat{a}\over\sigma(a)}\right)^{2}$ (79)
@@ B.16 The linear sigma in the log space
Equation (79) works for any positive, monotonicfunction $\sigma(a)$ as long as $\sigma(\hat{a}-{\sigma^{-}})={\sigma^{-}}$ and $\sigma(\hat{a}+{\sigma^{+}})={\sigma^{+}}$. As the values of $\sigma$ are supposed to be positive, it seems natural to interpolate these values in the log space. However, simple linear interpolation of $\ln\sigma$ as a function of $a$ is not going to work: for large values of $|a|$ the exponent of a linear function will grow faster than $|a-\hat{a}|$, and the magnitude of $\ln L(a)$ will decrease, creating at least one additional unphysical extremum.
Instead, it can be useful to perform linear interpolation of $\ln\sigma$ as a function of some variable $u=Q(a)$, where $Q(a)$ is a monotonictransform that maps the range of $a$ from [,][$\infty$,$\infty$][ - ∞ , ∞ ] to some compact interval $[u_{\text{min}},u_{\text{max}}]$. Without loss of generality, this interval can be chosen to be $[0,1]$. In this case, $Q(a)$ becomes a cumulative density function of some distribution supported on [,][$\infty$,$\infty$][ - ∞ , ∞ ].
@@
$\begin{split}\ln\sigma(Q(\hat{a}-{\sigma^{-}}))&=\ln{\sigma^{-}},\\ \ln\sigma(Q(\hat{a}+{\sigma^{+}}))&=\ln{\sigma^{+}}.\\ \end{split}$ (87)
@@ B.17 The double cubic sigma in the log space
$\ln\sigma(a)=\begin{cases}\ln{\sigma^{-}}&\text{for }a\leq\hat{a}-{\sigma^{-}}\\ P_{3}^{-}(a)&\text{for }\hat{a}-{\sigma^{-}}\leq a\leq\hat{a}\\ P_{3}^{+}(a)&\text{for }\hat{a}\leq a\leq\hat{a}+{\sigma^{+}}\\ \ln{\sigma^{+}}&\text{for }\hat{a}+{\sigma^{+}}\leq a\end{cases}$ (88)
@@
$\left.\frac{d}{da}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}=0,$
$\left.\frac{d^{2}}{da^{2}}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}=0,$
@@
$\left.\frac{d}{da}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}=0,$
$\left.\frac{d^{2}}{da^{2}}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}=0.$
@@ B.21 The skew normal
$\ln L(a)=-{1\over 2}\left({a-\xi\over\omega}\right)^{2}+\ln{\Phi\left(\alpha\left[{a-\xi)\over\omega}\right]\right)}$ (92)
@@ Appendix D Implementation in R
    returns the model parameters for a ’flipped’ OPAT result. direction is $+$1 (the default) when both deviations are positive and --1 when both are negative (the given values of both sp and sm are always positive. Only implemented for the dimidiated model.
@@
We give some simple illustrative examples. They are very basic, just to illustrate simple uses, and do not contain the checks and elaboration required for ‘good practice’. They assume that the library has been installed on your system and loaded and attached in this R session by library(AsymmetricErrors)
@@
      To combine two results for the Higgs width, using the linear variance model.
@@
      To plot the lnL curves for the result ${5.3}^{+1.2}_{-0.9}$ using all available models.
@@
      The near end of an object is ${100}^{+3}_{-4}$ km away, and the far end is ${200}^{+4}_{-5}$ km away. To calculate the error on the length of the object.
@@
      The left hand plot curves of Figure 4 were produced by the code
@@
      and the right hand plot is the same, with sp=0.85, sm=1.15 in place of sp=1.1, sm=0.9
@@
      If a table of OPAT-style systematic errors has been saved on a file errors.txt as bare numbers, one set per line, for ${\sigma^{+}}$ and ${\sigma^{-}}$, then to find the total, using the dimidiated model
@@
      To find the first three moments for a pdf specified as ${5.0}^{+1.1}_{-0.9}$ using all available models. (The substring call is to remove the Pdf. at the start of each name.)
@@
      To combine two measurements ${12.34}^{+0.56}_{-0.78}$ and ${12.43}^{+0.65}_{-0.87}$ using the railway Gaussian.
@@
      The combined results of Example 5.3 were produced by the code

### v3

@@ A.8 The Edgeworth expansion
The cumulative equivalent of this $p(x)$, the cumulative distribution function, is
@@
where $\Phi^{-1}$ is the inverse of the Gaussian cdf. This seems to work better than Newton’s method, thanks to the constantly changing gradient. From $z$ one obtains $M=\sigma z+\mu$, and thence ${\sigma^{+}}=\sigma+\mu-M$ and ${\sigma^{-}}=\sigma-\mu+M$.
@@ A.9 The skew normal
For the quantile parameters we need the cumulative distribution, which is not so elegant,
@@ A.10 The Johnson system
The combination of Johnson distributions $S_{B}$ (bounded), $S_{U}$ (unbounded), together with their limiting cases (Gaussian and log normal), form a very flexible system of four-parameter densities [23, 24, 25]. It turns out that there is a Johnson distribution for every possible combination of mean, standard deviation, skewness, and kurtosis. It is in fact convenient to think of the set as if it was a single distribution whose density is given by $j(R,\mu,\sigma,s,\kappa)$. (In this parameterization, skewness $s$ and kurtosis $\kappa$ are normalized.)
@@
Figure 26: The Johnson distribution. The panel on the right illustrates the Johnson $S_{U}$ density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Equation 48. As expected, this transformation maps 1-1- 1 to 0.6-0.6- 0.6, 0 to 0, and 111 into 111.
@@ A.11 The log normal
Figure 27: The Log Normal distribution. The panel on the right illustrates the density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the matching coordinate transformation calculated according to Equation 50.
@@ B.1 The cubic
If ${\sigma^{+}}$ and ${\sigma^{-}}$ differ by less than a factor of 2, this gives curves which behave sensibly in the range $[\hat{a}-{\sigma^{-}},\hat{a}+{\sigma^{+}}]$, but outside that range the cubic term gives an unwanted turning point. (If the asymmetry is large, the curve will have a minimum inside the range in addition to the maximum at $\hat{a}=0$.) Also the curve does not go to $\infty$- ∞ at both large positive and negative $a$. So although the cubic is the obvious extension of the quadratic, it often gives problems.
@@ B.4 The constrained quartic
$(\ln L)^{\prime\prime}(a)=-{1\over 2}(\alpha+\beta a)^{2}\qquad{\rm which\ yields}\qquad\ln L(a)=-{1\over 2}\left({\alpha^{2}a^{2}\over 2}+{\alpha\beta a^{3}\over 3}+{\beta^{2}a^{4}\over 12}\right).$ (56)
@@
$\beta={1\over{\sigma^{+}}{\sigma^{-}}}\sqrt{{12({\sigma^{-}}+{\sigma^{+}})^{2}\pm 24\sqrt{4{\sigma^{+}}{\sigma^{-}}^{3}+4{\sigma^{-}}{\sigma^{+}}^{3}-2{\sigma^{+}}^{4}-2{\sigma^{-}}^{4}}\over 3{\sigma^{-}}^{2}+2{\sigma^{+}}{\sigma^{-}}+3{\sigma^{+}}^{2}}}.$ (57)
@@ B.7 The interpolated 7th degree polynomial
$\begin{split}\alpha&=6({\sigma^{-}}-{\sigma^{+}})/\eta\\ \beta&=15({\sigma^{-}}-{\sigma^{+}})^{2}/\eta\\ \gamma&=10({\sigma^{-}}-{\sigma^{+}})({\sigma^{-}}^{2}-4{\sigma^{-}}{\sigma^{+}}+{\sigma^{+}}^{2})/\eta\\ \delta&=-30{\sigma^{-}}{\sigma^{+}}({\sigma^{-}}-{\sigma^{+}})^{2}/\eta\\ \epsilon&=30{\sigma^{-}}^{2}{\sigma^{+}}^{2}({\sigma^{-}}-{\sigma^{+}})/\eta\\ \zeta&=({\sigma^{-}}^{6}+4{\sigma^{-}}^{5}{\sigma^{+}}+6{\sigma^{-}}^{4}{\sigma^{+}}^{2}-6{\sigma^{-}}^{3}{\sigma^{+}}^{3}+6{\sigma^{-}}^{2}{\sigma^{+}}^{4}+4{\sigma^{-}}{\sigma^{+}}^{5}+{\sigma^{+}}^{6})/\eta\end{split}.$ (64)
@@
[…]
@@ B.9 The double quintic
$\ln L(a)=\begin{cases}P_{2}^{-}(a)&\text{for }a\leq-{\sigma^{-}}\\ P_{5}^{-}(a)&\text{for }-{\sigma^{-}}\leq a\leq 0\\ P_{5}^{+}(a)&\text{for }0\leq a\leq{\sigma^{+}}\\ P_{2}^{+}(a)&\text{for }{\sigma^{+}}\leq a\end{cases},$ (68)
@@
The simple double quintic is obtained by setting $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$. For this choice of $\sigma_{0}$, the requirement that $\ln L(a)$ has a single extremum (the maximum at 0) leads to the condition that ${\sigma^{+}}$ and ${\sigma^{-}}$ must differ by less than a factor of 13.513.513.5.
@@ B.11 The log logistic-beta
[…]
@@
Equation 72 satisfies the conditions $\ln L(0)=0$ and $\left.\frac{d\ln L(a)}{da}\right|_{a=0}=0$ automatically. It also needs to satisfy the conditions $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})=-\frac{1}{2}$. While with three parameters $A$, $c$, and $g$ the latter conditions can be satisfied in a multitude of ways, we choose the parameterization that minimizes $g$, i.e., $A$ and $c$ are chosen in such a way that $|\ln L({\sigma^{+}})/g|$ is maximized subject to the constraint $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})$. This results in a conservative log-likelihood model with limited slopes and the second derivative exceeding $\frac{1}{\min({\sigma^{-}},{\sigma^{+}})^{2}}$ in magnitude only on a short interval.
@@ B.13 The generalised Poisson
This has to be solved numerically for $\gamma$. The solution lies between 0 and $1/{\sigma^{-}}$, and can be found by repeatedly taking the midpoint [^5] Attempts to use more sophisticated algorithms were unsuccessful.. ${\cal N}$ is then found from
@@
This does fairly well, though the need for a numerical solution makes it inelegant. If the curve to be fitted has a negative skewness (${\sigma^{-}}>{\sigma^{+}})$, then the sign of $a$ has to be flipped.
@@ B.14 The linear sigma
$\ln L(a)=-{1\over 2}\left({a-\hat{a}\over\sigma(a)}\right)^{2}.$ (79)
@@ B.16 The linear sigma in the log space
[…]
Instead, it can be useful to perform linear interpolation of $\ln\sigma$ as a function of some variable $u=Q(a)$, where $Q(a)$ is a monotonic transform that maps the range of $a$ from [,][$\infty$,$\infty$][ - ∞ , ∞ ] to some compact interval $[u_{\text{min}},u_{\text{max}}]$. Without loss of generality, this interval can be chosen to be [0,1][0,1][ 0 , 1 ]. In this case, $Q(a)$ becomes a cumulative density function of some distribution supported on [,][$\infty$,$\infty$][ - ∞ , ∞ ].
@@
$\begin{split}\ln\sigma(Q(\hat{a}-{\sigma^{-}}))&=\ln{\sigma^{-}}\\ \ln\sigma(Q(\hat{a}+{\sigma^{+}}))&=\ln{\sigma^{+}}.\\ \end{split}$ (87)
@@ B.17 The double cubic sigma in the log space
$\ln\sigma(a)=\begin{cases}\ln{\sigma^{-}}&\text{for }a\leq\hat{a}-{\sigma^{-}}\\ P_{3}^{-}(a)&\text{for }\hat{a}-{\sigma^{-}}\leq a\leq\hat{a}\\ P_{3}^{+}(a)&\text{for }\hat{a}\leq a\leq\hat{a}+{\sigma^{+}}\\ \ln{\sigma^{+}}&\text{for }\hat{a}+{\sigma^{+}}\leq a\end{cases},$ (88)
@@
$\left.\frac{d}{da}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}$ $=$ 0 0,0 ,
$\left.\frac{d^{2}}{da^{2}}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}$ $=$ 0 0,0 ,
@@
$\left.\frac{d}{da}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}$ $=$ 0 0,0 ,
$\left.\frac{d^{2}}{da^{2}}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}$ $=$ 0 0.0 .
@@ B.21 The skew normal
$\ln L(a)=-{1\over 2}\left({a-\xi\over\omega}\right)^{2}+\ln{\Phi\left(\alpha\left[{a-\xi)\over\omega}\right]\right)}.$ (92)
@@ Appendix D Implementation in R
    returns the model parameters for a ‘flipped’ OPAT result. direction is $+$1 (the default) when both deviations are positive and --1 when both are negative (the given values of both sp and sm are always positive). Only implemented for the dimidiated model.
@@
We give some simple illustrative examples. They are very basic, just to illustrate simple uses, and do not contain the checks and elaboration required for ‘good practice’. They assume that the library has been installed on your system and loaded and attached in this R session by library(AsymmetricErrors).
@@
      To combine two results for the Higgs width, using the linear variance model:
@@
      To plot the lnL curves for the result ${5.3}^{+1.2}_{-0.9}$ using all available models:
@@
      The near end of an object is ${100}^{+3}_{-4}$ km away, and the far end is ${200}^{+4}_{-5}$ km away. To calculate the error on the length of the object:
@@
      The left hand plot curves of Figure 4 were produced by the code:
@@
      and the right hand plot is the same, with sp=0.85, sm=1.15 in place of sp=1.1, sm=0.9.
@@
      If a table of OPAT-style systematic errors has been saved on a file errors.txt as bare numbers, one set per line, for ${\sigma^{+}}$ and ${\sigma^{-}}$, then to find the total, using the dimidiated model:
@@
      To find the first three moments for a pdf specified as ${5.0}^{+1.1}_{-0.9}$ using all available models:
@@
      (The substring call is to remove the Pdf. at the start of each name.)
@@
      To combine two measurements ${12.34}^{+0.56}_{-0.78}$ and ${12.43}^{+0.65}_{-0.87}$ using the railway Gaussian:
@@
      The combined results of Example 5.3 were produced by the code: